OpenStudy (anonymous):

Solve the equation for x: 10x = 27.17 Choose one answer. a. -0.66 b. -0.43 c. 1.43 d. 1.83

OpenStudy (anonymous):

a. -0.66

OpenStudy (anonymous):

Solve log5 (x + 1) - log5 (x - 2) = log54 for x. Choose one answer. a. -2 b. 1/3 c. 1 d. 3

OpenStudy (anonymous):

c. 1 may i ask if this is homework or a test

OpenStudy (anonymous):

quiz

OpenStudy (anonymous):

k then i guess i can help you

OpenStudy (anonymous):

ok thanks cuz i really need help Solve log4 8 - log4 (x + 2) = log44 for x. Choose one answer. a. 0 b. 1.5 c. 2.5 d. 4

OpenStudy (anonymous):

try b.1.5 im sure thats correct

OpenStudy (anonymous):

okSolve the equation for x: x = log111331 Choose one answer. a. -1/3 b. 1/3 c. 3 d. 4 now wouldnt this one be -1/3 or just 3

OpenStudy (anonymous):

8========================D APPLE POO

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Solve the equation for x: -2 = logx(1/144). Choose one answer. a. 1/13 b. 1/12 c. 12 d. 13 now this one is b right

OpenStudy (anonymous):

man im jking i have no idea wat im evan talking about srry

OpenStudy (anonymous):

so u dont know this

OpenStudy (anonymous):

i no that aint no quiz u doing a test

OpenStudy (anonymous):

testttttttttttttttt

OpenStudy (anonymous):

its a quiz its only 10 question